the theoretical yield of 2cuo+c?2cu+co2

s(a 0zt ODM 001 5 m .G mol Test bbe 314? wu 3 36 ,33 , 34 32 2.0 ot /2016

we used 2.04 g of CuO and .654 g of C
Answer #1

2Cuo + C -----> 2Cu +CO2

Theortical mass of Cu produced Cuo = 2.04 gms x 2 moles of Cux 63.54 mol wt of Cu /79.54 mol wt of CuO X 2 moles of CuO = 259.24/159.08 = 1.629 gms

Theortical mass of Cu produced C = 0.654 gms x 2 moles of Cu x 63.54 mol wt of Cu /14 mol wt of C X 1 moles of C = 83.11/14 = 5.936 gms

The one which produces less amount of theoritical mass of product is limting reagent so CuO (cuprous oxide) is the limting reagent and theortical yield is 1.629 gms of Cu(copper)